### v2

@@ Appendix D Single-Action Auctions (Section 5) Appendix
For any single-action, optimal auction, there exist unique $\tilde{x}^{*}:\mathcal{V}^{N}\to\{0,1\}^{N}$ and $\tilde{t}:\mathcal{V}^{N}\to\mathbb{R}^{N}$ such that:
@@
    (Correspondence) For all $v\in\mathcal{V}^{N}$, $\tilde{x}^{*}(v)=x(\sigma(v;B))$ and $\tilde{t}(v)=t\left(\sigma(v;B)\right)$.
@@
    (Incentive Compatibility) For all $R$, $i\in R$, $v_{i}$, and $v_{i}^{\prime}$,
    $\mathds{E}_{v_{-i},\tilde{R}}\left[\tilde{x}^{*}_{i}\left(\sigma(v;\tilde{R})\right)v_{i}-\tilde{t}_{i}(\sigma(v;\tilde{R}))\mid v_{i},s_{i}=\psi(v_{j<i})\right]\\ \geq\mathds{E}_{v_{-i},\tilde{R}}\left[\tilde{x}^{*}_{i}\left(\sigma(v_{i}^{\prime},v_{-i};\tilde{R})\right)v_{i}-\tilde{t}_{i}\left(\sigma(v_{i}^{\prime},v_{-i};\tilde{R})\right)\mid v_{i},s_{i}=\psi(v_{j<i})\right].$ (9)
Proof.
To begin, let us note that we can uniquely define $(\tilde{x}^{*},\tilde{t})$ point-wise from $\sigma(v;B)$. The direct allocation rule is uniquely defined as $\tilde{x}^{*}$ for all optimal auctions. Next, the IR constraint (Condition ii) follows immediately from the ex-post IR condition and our construction of $(\tilde{x}^{*},\tilde{t})$. Finally, Equation 9 comes from Definition A.1 and recalling that we restrict shill bidders to actions that could have been taken by real bidders. ∎
Lemma D.2.
@@
$\mathds{E}_{v}\left[\sum_{k\in R}\tilde{t}_{k}\left(\left\{v_{i}\right\}_{i\in S},\left\{v_{i}\right\}_{i\notin S}\right)\mid v_{j<\min S}\right]\leq\mathds{E}_{v}\left[\sum_{k\in R}\tilde{t}_{k}\left(0,\left\{v_{i}\right\}_{i\notin S}\right)\mid v_{j<\min S}\right].$
Proof.
Towards contradiction, suppose there exists $R$, $v_{j<\min S}$, and $\left\{v_{i}\right\}_{i\in S}$ such that
$\mathds{E}_{v}\left[\sum_{k\in R}\tilde{t}_{k}\left(\left\{v_{i}\right\}_{i\in S},\left\{v_{i}\right\}_{i\notin S}\right)\mid v_{j<\min S}\right]>\mathds{E}_{v}\left[\sum_{k\in R}\tilde{t}_{k}\left(0,\left\{v_{i}\right\}_{i\notin S}\right)\mid v_{j<\min S}\right].$
We now prove that the deviation by the coalition $S$ where they report $\left\{v_{i}\right\}_{i\in S}$ is profitable and therefore that the auction is not weakly shill-proof. By assumption, a shill bidder observes actions by all bidders who take actions before her. So, $\left\{v_{i}\right\}_{i\in S}$ can condition on $v_{j<\min S}$ when making decisions. Then, the strategy by $S$ of committing to report $\left\{v_{i}\right\}_{i\in S}$ regardless of what other bidders play after $\min_{i\in S}\left\{i\right\}$ must be strictly profitable compared to always reporting 00. Thus, we have found a strategy that does strictly better than always reporting 00: When the values before $\min_{i\in S}\left\{i\right\}$ are reported as $v_{j<\min S}$, report $\left\{v_{i}\right\}_{i\in S}$. Otherwise, report 00. This strategy in the direct game immediately translates to a profitable deviation in the auction by Definition A.1 and Lemma D.1 and thus the equilibrium is not weakly shill-proof. ∎
Lemma D.3.
If a single-action, optimal auction is strongly shill-proof, then for all $R$, $i\notin R$, $v_{i}$, and $v_{-i}$, $\sum_{k\in R}\tilde{t}_{k}(v_{i},v_{-i})\leq\sum_{k\in R}\tilde{t}_{k}(0,v_{-i}).$
Proof.
Towards contradiction, suppose that there exists $R$, $i\notin R$, $v_{i}$, and $v_{-i}$ such that $\sum_{k\in R}\tilde{t}_{k}(v_{i},v_{-i})>\sum_{k\in R}\tilde{t}_{k}(0,v_{-i}).$ That means in the direct game, reporting 00 is not a dominant strategy for shill bidders. This implies, from Condition i of Lemma D.1, that there exists a deviation in the auction such that for some value vectors, the seller raises more revenue. Therefore, the auction is not strongly shill-proof. ∎
Now, when discussing single-action, optimal auctions, we focus on the direct mechanisms associated to weakly shill-proof auctions and so we will refer to an auction as $(\tilde{x},\tilde{t},\psi)$ without reference to $R$.
Lemma D.4.
Suppose a single-action, optimal auction $(\tilde{x}^{*},\tilde{t},\psi)$ is weakly shill-proof. Then, for all $i$, $v$, and $v_{j>i}^{\prime}$, $\left[\tilde{x}^{*}_{i}\left(v\right)=\tilde{x}^{*}_{i}\left(v_{j\leq i},v_{j>i}^{\prime}\right)\implies\tilde{t}_{i}\left(v\right)=\tilde{t}_{i}\left(v_{j\leq i},v_{j>i}^{\prime}\right)\right].$
Proof.
Towards contradiction, suppose there exists $i$, $v$, and $v_{j>i}^{\prime}$, such that $\tilde{x}^{*}_{i}\left(v\right)=\tilde{x}^{*}_{i}\left(v_{j\leq i},v_{j>i}^{\prime}\right)$, but $\tilde{t}_{i}\left(v;s\right)>\tilde{t}_{i}\left(v_{j\leq i},v_{j>i}^{\prime};s\right)$. Because the auction is winner-paying, it must then be the case that $\tilde{x}^{*}_{i}\left(v\right)=\tilde{x}^{*}_{i}\left(v_{j\leq i},v_{j>i}^{\prime}\right)=1$. Let $R=\left\{1,\ldots,i\right\}$. Then,
$\mathds{E}_{v}\left[\sum_{k\in R}\tilde{t}_{k}(\left\{v_{i}\right\}_{i\in S},\left\{v_{i}\right\}_{i\notin S})\mid v_{j\leq\min S}\right]$ $=\tilde{t}_{i}\left(v\right)>\tilde{t}_{i}\left(v_{j\leq i},v_{j>i}^{\prime}\right)\geq\tilde{t}_{i}\left(v_{j\leq i},0\right).$
This violates Lemma D.2, and so we have reached a contradiction. ∎
Lemma D.5.
Suppose a single-action, optimal auction $(\tilde{x}^{*},\tilde{t},\psi)$ is mildly ex-post incentive compatible and weakly shill-proof. Then, there exists $i<N$ such that for all $v_{i}$, $v_{i}^{\prime}$, and $v_{-i}\in\psi^{-1}_{i}(v_{j<i})$, $\left[\tilde{x}^{*}_{i}(v)=\tilde{x}^{*}_{i}(v_{i}^{\prime},v_{-i})\implies\tilde{t}_{i}(v)=\tilde{t}_{i}(v_{i}^{\prime},v_{-i})\right].$
Proof.
Let $i<N,R\ni i,v_{i},v_{i}^{\prime},v_{-i}\in\psi^{-1}_{i}(v_{j<i})$ such that $\tilde{x}^{*}_{i}(v)=\tilde{x}^{*}_{i}(v_{i}^{\prime},v_{-i})$. WLOG, suppose $v_{i}>v_{i}^{\prime}$. By monotonicity, $\tilde{t}_{i}(v)\geq\tilde{t}_{i}(v_{i}^{\prime},v_{-i})$. Towards contradiction, suppose $\tilde{t}_{i}(v)>\tilde{t}_{i}(v_{i}^{\prime},v_{-i})$. By the winner-paying property, $\tilde{t}_{i}(v)>\tilde{t}_{i}(v_{i}^{\prime},v_{-i})$ implies that $\tilde{x}^{*}_{i}(v)=\tilde{x}^{*}_{i}(v_{i}^{\prime},v_{-i})=1$. However, note that $\tilde{t}_{i}(v)>\tilde{t}_{i}(v_{i}^{\prime},v_{-i})$ would mean that the utility of reporting $v_{i}^{\prime}$ would be higher than truthful reporting under true value $v_{i}$ which would violate the mildly ex-post incentive compatiblility and thus $\tilde{t}_{i}(v_{i}^{\prime},v_{-i})=\tilde{t}_{i}(v_{i}^{\prime},v_{-i})$. ∎
Proof of Theorem 5.2.
Towards contradiction, suppose such an auction did exist. Fix $i<N,s$ and suppose $v_{i}<v^{M}$. Combining Lemmatta D.4 and D.5, we can see that for all $v_{i}^{\prime}$ and $v_{-i},v_{-i}^{\prime}\in\psi^{-1}_{i}(v_{j<i})$, $\left[\tilde{x}^{*}_{i}(v)=\tilde{x}^{*}_{i}(v^{\prime})\implies\tilde{t}_{i}(v)=\tilde{t}_{i}(v^{\prime})\right].$ So, define $\tilde{t}_{i}^{*}$ as the (constant) $\tilde{t}_{i}(v)$ for all $v$ such that $\tilde{x}^{*}_{i}(v)=1$.
By definition, when $\tilde{x}^{*}_{i}(v)=1$, it must also be the case that $\tilde{x}^{*}_{i}(v^{M},v_{-i})=1$. So, applying the winner-paying property (and suppressing that the expectation is conditioned on $s_{i}=\psi_{i}(v_{j<i})$), we have
$\mathds{E}_{v_{-i}}\left[\tilde{x}^{*}_{i}\left(v^{M},v_{-i}\right)v_{i}-\tilde{t}_{i}\left(v^{M},v_{-i}\right)\right]\\ =\mathds{E}_{v_{-i}}\left[v_{i}-\tilde{t}_{i}^{*}\mid\tilde{x}^{*}_{i}(v)=1\right]+\mathds{E}_{v_{-i}}\left[v_{i}-\tilde{t}_{i}^{*}\mid\tilde{x}^{*}_{i}(v)=0,\tilde{x}^{*}_{i}\left(v^{M},v_{-i}\right)=1\right],$ (10)
and $\text{and }\mathds{E}_{v_{-i}}\left[\tilde{x}^{*}_{i}\left(v\right)v_{i}-\tilde{t}_{i}\left(v\right)\right]=\mathds{E}_{v_{-i}}\left[v_{i}-\tilde{t}_{i}^{*}\mid\tilde{x}^{*}_{i}(v)=1\right].$ (11)
Taking the difference between Equation 10 and Equation 11, we see that
$\mathds{E}_{v_{-i}}\left[\tilde{x}^{*}_{i}\left(v^{M},v_{-i}\right)v_{i}-\tilde{t}_{i}\left(v^{M},v_{-i}\right)\right]-\mathds{E}_{v_{-i}}\left[\tilde{x}^{*}_{i}\left(v\right)v_{i}-\tilde{t}_{i}\left(v\right)\right]\\ =\mathds{E}_{v_{-i}}\left[v_{i}-\tilde{t}_{i}^{*}\mid\tilde{x}^{*}_{i}(v)=0,\tilde{x}^{*}_{i}\left(v^{M},v_{-i}\right)=1\right]$ (12)
Now, by definition $\tilde{x}^{*}$ is monotone, and by assumption $\tilde{t}$ is monotone. If there exists $v^{m}$ such that $\mathds{P}[\tilde{x}^{*}_{i}(v^{m},v_{-i})=1]<\mathds{P}[\tilde{x}^{*}_{i}(v^{M},v_{-i})=1]$, then for such $m$,
$\mathds{E}_{v_{-i}}\left[\tilde{x}^{*}_{i}(v^{m},v_{-i})v^{m}-\tilde{t}_{i}(v^{m},v_{-i})\right]\geq\mathds{E}_{v_{-i}}\left[\tilde{x}^{*}_{i}(v^{m-1},v_{-i})v^{m}-\tilde{t}_{i}(v^{m-1},v_{-i})\right]$
$>\mathds{E}_{v_{-i}}\left[\tilde{x}^{*}_{i}(v^{m-1},v_{-i})v^{m-1}-\tilde{t}_{i}(v^{m-1},v_{-i})\right]\geq 0,$
where the last inequality comes from the IR condition. Thus, the IR constraint does not bind for $v_{i}=v^{m}$. Since the good has to be allocated to the highest type, for all $i<N$, there exists $v_{-i}$ such that $\tilde{x}^{*}_{i}(v)=0$ and $\tilde{x}^{*}_{i}(v^{M},v_{-i})=1$. Thus,
$\mathds{E}_{v_{-i}}\left[v_{i}-\tilde{t}_{i}^{*}\mid\tilde{x}^{*}_{i}(v)=0,\tilde{x}^{*}_{i}\left(v^{M},v_{-i}\right)=1\right]>0.$ (13)
Combining Equations 12 and 13, we see that
$\mathds{E}_{v_{-i}}\left[\tilde{x}^{*}_{i}\left(v^{M},v_{-i}\right)v_{i}-\tilde{t}_{i}\left(v^{M},v_{-i}\right)\right]>\mathds{E}_{v_{-i}}\left[\tilde{x}^{*}_{i}\left(v\right)v_{i}-\tilde{t}_{i}\left(v\right)\right].$ (14)
@@
$\mathds{E}_{v_{-i}}\left[\tilde{x}^{*}_{i}\left(v\right)v_{i}-\tilde{t}_{i}(v)\right]\geq\mathds{E}_{v_{-i}}\left[\tilde{x}^{*}_{i}\left(v_{i}^{\prime},v_{-i}\right)v_{i}-\tilde{t}_{i}\left(v_{i}^{\prime},v_{-i}\right)\right].$
Taking $v_{i}^{\prime}=v^{M}$, Equation 14 violates the IC constraint from Lemma D.1—and thus we have reached a contradiction. ∎
Appendix O Online Appendix

### v3

@@ Appendix D Affiliation and Interdependence (Section 5) Appendix
For any function $g:\mathbb{R}^{N}\to\mathbb{R}$ with decreasing differences, if $F^{\prime}\succeq_{\mathrm{Aff}}F$, then for any index $i$ and for all $\theta^{1}_{i}>\theta^{2}_{i}>\theta^{3}_{i}$,
$\mathds{E}_{\theta\sim F^{\prime}}$ $\left[\left(g(\theta^{1}_{i},\theta_{-i})-g(\theta^{2}_{i},\theta_{-i})\right)-\left(g(\theta^{2}_{i},\theta_{-i})-g(\theta^{3}_{i},\theta_{-i})\right)\right]$
$\leq\mathds{E}_{\theta\sim F}\left[\left(g(\theta^{1}_{i},\theta_{-i})-g(\theta^{2}_{i},\theta_{-i})\right)-\left(g(\theta^{2}_{i},\theta_{-i})-g(\theta^{3}_{i},\theta_{-i})\right)\right].$
Proof.
Consider any $i$, $\Delta^{1}$, and $\Delta^{2}$. Then, define
$\tilde{g}(\theta_{i},\theta_{-i})=\left(g(\theta_{i}+\Delta^{1}+\Delta^{2},\theta_{-i})-g(\theta_{i}+\Delta^{1},\theta_{-i})\right)-\left(g(\theta_{i}+\Delta^{1},\theta_{-i})-g(\theta_{i},\theta_{-i})\right).$
By the assumption that $g$ has decreasing differences, $\tilde{g}$ is non-increasing in $\theta_{i}$. We can then immediately apply Lemma D.1 to complete the proof. ∎
Proof of Lemma 5.5
To prove that $T$ has the claimed form, we will consider a specific non-truthful reporting: if a bidder has value $\theta^{m}$, she commits to mis-reporting (selecting partitions) $\theta^{m^{\prime}}$ for the rest of the game. Since our direct mechanism is an equilibrium for real bidders, we must have
$\mathds{E}\Big[\Big(X_{i}(\theta^{j_{m}};$ $\Theta)v(\theta^{j_{m}},\theta_{-i})-T_{i}(\theta^{j_{m}};\Theta)\Big)$-
$\left(X_{i}(\theta^{j_{m-1}};\Theta)v(\theta^{j_{m}},\theta_{-i})-T_{i}(\theta^{j_{m-1}};\Theta)\right)\mid\theta_{i}=\theta^{j_{m}},\theta_{-i}\in\Theta_{-i}\Big]\geq 0,\text{ and}$
$\mathds{E}\Big[\Big(X_{i}(\theta^{j_{m-1}};$ $\Theta)v(\theta^{j_{m-1}},\theta_{-i})-T_{i}(\theta^{j_{m-1}};\Theta)\Big)$-
$\left(X_{i}(\theta^{j_{m}};\Theta)v(\theta^{j_{m-1}},\theta_{-i})-T_{i}(\theta^{j_{m}};\Theta)\right)\mid\theta_{i}=\theta^{j_{m-1}},\theta_{-i}\in\Theta_{-i}\Big]\geq 0.$
$U_{i}(\theta^{j_{m}};\Theta)\geq U_{i}(\theta^{j_{m-1}};\Theta)+(v^{*}_{i}(\theta^{j_{m-1}},\theta^{j_{m}};\Theta)-v^{*}_{i}(\theta^{j_{m-1}},\theta^{j_{m-1}};\Theta))X_{i}(\theta^{j_{m-1}};\Theta),\text{ and}$
$U_{i}(\theta^{j_{m-1}};\Theta)\geq U_{i}(\theta^{j_{m}};\Theta)+(v^{*}_{i}(\theta^{j_{m}},\theta^{j_{m}};\Theta)-v^{*}_{i}(\theta^{j_{m}},\theta^{j_{m-1}};\Theta))X_{i}(\theta^{j_{m}};\Theta).$
Thus,
$(v^{*}_{i}(\theta^{j_{m-1}},$ $\theta^{j_{m}};\Theta)-v^{*}_{i}(\theta^{j_{m-1}},\theta^{j_{m-1}};\Theta))X_{i}(\theta^{j_{m-1}};\Theta)$
$\leq U_{i}(\theta^{j_{m}};\Theta)-U_{i}(\theta^{j_{m-1}};\Theta)$
$\leq(v^{*}_{i}(\theta^{j_{m}},\theta^{j_{m}};\Theta)-v^{*}_{i}(\theta^{j_{m}},\theta^{j_{m-1}};\Theta))X_{i}(\theta^{j_{m}};\Theta).$
Then, by the same logic as in Lemma B.2, a mechanism with an optimal transfer rule is such that
$U_{i}(\theta^{j_{m}};\Theta)=\sum_{k=2}^{m}(v^{*}_{i}(\theta^{j_{k-1}},\theta^{j_{k}};\Theta)-v^{*}_{i}(\theta^{j_{k-1}},\theta^{j_{k-1}};\Theta))X_{i}(\theta^{j_{k-1}};\Theta).$
Therefore, we have that
$T_{i}(\theta^{j_{m}};\Theta)=X_{i}(\theta^{j_{m}};\Theta)v^{*}_{i}(\theta^{j_{m}},\theta^{j_{m}};\Theta)-\sum_{k=2}^{m}(v^{*}_{i}(\theta^{j_{k-1}},\theta^{j_{k}};\Theta)-v^{*}_{i}(\theta^{j_{k-1}},\theta^{j_{k-1}};\Theta))X_{i}(\theta^{j_{k-1}};\Theta).$ (15)
Re-arranging concludes the proof. ∎
Proof of Lemma 5.6
There are two cases to consider. In the first case, suppose $\max_{\theta\in S}\left\{\theta\right\}>\max_{\theta\in S^{\prime}}\left\{\theta\right\}$. In this case, we can form subsets $S_{1},\ldots,S_{K}$ and $S^{\prime}_{1},\ldots,S^{\prime}_{K}$ such that
@@
    $S=\bigcup_{k}S_{k}$ and $S^{\prime}=\bigcup_{k}S^{\prime}_{k}$;
@@
    For all $k^{\prime}>k$, $x\in S_{k}$, and $y\in S^{\prime}_{k^{\prime}}$: $x<y$.
Such partitions can be formed inductively. First, define the base case as
@@
Then, define the inductive case as
[…]
$S^{\prime}_{k}=\left\{\theta\in S^{\prime}\ \backslash\ \left(\bigcup_{k^{\prime}>k}S_{k^{\prime}}\right):\theta>\max_{\tilde{\theta}\in S\backslash\left(\bigcup_{k^{\prime}\geq k}S_{k^{\prime}}\right)}\left\{\tilde{\theta}\right\}\right\}$.
Note that these two partitions have the same number of elements because $\min_{\theta\in S}\left\{\theta\right\}>\min_{\theta\in S^{\prime}}\left\{\theta\right\}$. Then, we can re-write the right-hand side of Equation 1 as
$\sum_{k=1}^{K}\left(\mathds{P}_{F}\left[\theta_{i}\in S_{k}\mid S\right]\mathds{E}_{\theta\sim F}\left[g(\theta)\mid\theta_{i}\in S_{k}\right]-\mathds{P}_{F}\left[\theta_{i}\in S^{\prime}_{k}\mid S^{\prime}\right]\mathds{E}_{\theta\sim F}\left[g(\theta)\mid\theta_{i}\in S^{\prime}_{k}\right]\right).$
Note that by assumption, the marginals are equal: $\mathds{P}_{F^{\prime}}\left[\theta_{1}\in S_{k}\right]=\mathds{P}_{F}\left[\theta_{1}\in S_{k}\right]$ for all $k$. We can then recall that $g$ is non-decreasing, weakly super-modular, and that affiliation is equivalent to log-supermodularity of the type distribution to conclude that for all $k$,
$\mathds{P}_{F^{\prime}}$ $\left[\theta_{i}\in S_{k}\mid S\right]\mathds{E}_{\theta\sim F^{\prime}}\left[g(\theta)\mid\theta_{i}\in S_{k}\right]-\mathds{P}_{F^{\prime}}\left[\theta_{i}\in S^{\prime}_{k}\mid S^{\prime}\right]\mathds{E}_{\theta\sim F^{\prime}}\left[g(\theta)\mid\theta_{i}\in S^{\prime}_{k}\right]$
$\geq\mathds{P}_{F}\left[\theta_{i}\in S_{k}\mid S\right]\mathds{E}_{\theta\sim F}\left[g(\theta)\mid\theta_{i}\in S_{k}\right]-\mathds{P}_{F}\left[\theta_{i}\in S^{\prime}_{k}\mid S^{\prime}\right]\mathds{E}_{\theta\sim F}\left[g(\theta)\mid\theta_{i}\in S^{\prime}_{k}\right]$
Then, we can sum over all $k$ to conclude the proof in this case.
Now, we consider the case where $\max_{\theta\in S}\left\{\theta\right\}<\max_{\theta\in S^{\prime}}\left\{\theta\right\}$. In this case we can form subsets $S_{1},\ldots,S_{K}$ and $S^{\prime}_{1},\ldots,S^{\prime}_{K+1}$ such that
  1. (i)
    $S=\bigcup_{k}S_{k}$ and $S^{\prime}=\bigcup_{k}S^{\prime}_{k}$;
  2. (ii)
    For all $k$, $x\in S_{k}$, and $y\in S^{\prime}_{k}$: $x>y$; and
  3. (iii)
    For all $k^{\prime}>k$, $x\in S_{k}$, and $y\in S^{\prime}_{k^{\prime}}$: $x<y$.
The inductive construction is the same as the first case with the base case changed to
[…]
We then re-write the right-hand side of Equation 1 as
$\frac{1}{2}$ $\mathds{P}_{F}\left[\theta_{i}\in S_{1}\right]\mathds{E}_{\theta\sim F}\left[g(\theta)\mid\theta_{i}\in S_{1}\right]-\mathds{P}_{F}\left[\theta_{i}\in S^{\prime}_{1}\right]\mathds{E}_{\theta\sim F}\left[g(\theta)\mid\theta_{i}\in S^{\prime}_{1}\right]+$
$\frac{1}{2}$ $\mathds{P}_{F}\left[\theta_{i}\in S_{K}\right]\mathds{E}_{\theta\sim F}\left[g(\theta)\mid\theta_{i}\in S_{K}\right]-\mathds{P}_{F}\left[\theta_{i}\in S^{\prime}_{K+1}\right]\mathds{E}_{\theta\sim F}\left[g(\theta)\mid\theta_{i}\in S^{\prime}_{K+1}\right]+$
$\frac{1}{2}\sum_{k=2}^{K}\Bigg(\left(\mathds{P}_{F}\left[\theta_{i}\in S_{k}\right]\mathds{E}_{\theta\sim F}\left[g(\theta)\mid\theta_{i}\in S_{k}\right]-\mathds{P}_{F}\left[\theta_{i}\in S^{\prime}_{k}\right]\mathds{E}_{\theta\sim F}\left[g(\theta)\mid\theta_{i}\in S^{\prime}_{k}\right]\right)$
@@
$\qquad\quad-\left(\mathds{P}_{F}\left[\theta_{i}\in S^{\prime}_{k}\right]\mathds{E}_{\theta\sim F}\left[g(\theta)\mid\theta_{i}\in S^{\prime}_{k}\right]-\mathds{P}_{F}\left[\theta_{i}\in S_{k-1}\right]\mathds{E}_{\theta\sim F}\left[g(\theta)\mid\theta_{i}\in S_{k-1}\right]\right)\Bigg)$
Now, observe that the first two differences must decrease under $F^{\prime}$ because $g$ is non-decreasing and weakly super-modular. We can then recall that $g$ has decreasing differences, apply Lemma D.2, and then sum over all $k$ to conclude that the sum also must decrease and therefore obtain the desired conclusion. ∎
Proof of Theorem 5.4
Affiliation.
Towards contrapositive, suppose that the auction is not shill-proof (strong or weak, we will highlight where the proofs diverge) under $F$ and consider some $F^{\prime}\succeq_{\mathrm{Aff}}F$. There exists some realization of $\theta$ under $F$ such that a shill bidder $i$ deviates at $\Theta$. Since $\tilde{x}$, $\mu$, and $\xi_{0}$ are all fixed, we know from Lemma 5.5 that under weak shill-proofness that the only way for expected revenue under weak shill-proofness to change is for the ex-interim expected value of some bidder $j\neq i$ to change. We can also apply Lemma B.3 to see that the same is true under strong shill-proofness. Thus, in order to prove that the auction is not shill-proof for either weak or strong shill-proofness under $F^{\prime}$, it is sufficient to show that for all $j\neq i$, if the expected transfers from $T_{j}$ from manipulation $\hat{\Theta}$ compared to truthful play $\Theta^{*}$ under $F$, then the same manipulation is profitable under $F^{\prime}$. Fix $j$ and consider the profitably manipulated possible values $\hat{\Theta}$. We can re-express the expected transfers from bidder $j$ from Lemma 5.5 as
$\mathds{E}$ $_{\theta\sim F}\left[T_{j}(\theta_{j};\Theta^{*})\mid\theta\in\Theta^{*}\right$]
$=\mathds{E}_{\theta\sim F}\Bigg[\mathds{E}\left[\tilde{x}_{j}(\theta_{j},\theta_{-j})v(\theta_{j},\theta_{-j})\mid\theta_{j}=\theta_{j},\tilde{x}_{j}(\theta_{j},\theta_{-j})=1\right]$
$\qquad-\sum_{m:\theta^{k_{m}}<\theta_{j}}\mathds{P}\left[\tilde{x}_{j}(\theta^{k_{m}},\theta_{-j})=1\right]\big(\mathds{E}\left[v(\theta^{k_{m}},\theta_{-j})\mid\theta_{j}=\theta^{k_{m+1}},\tilde{x}_{j}(\theta^{k_{m}},\theta_{-j})=1\right]$
$\qquad\qquad\qquad\qquad-\mathds{E}\left[v(\theta^{k_{m}},\theta_{-j})\mid\theta_{j}=\theta^{k_{m}},\tilde{x}_{j}(\theta^{k_{m}},\theta_{-j})=1\right]\big)\mid\theta_{-j}\in\Theta_{-j}^{*},\theta_{j}\in\Theta^{*}_{j}\Bigg].$
Now observe that with manipulation only changes that the expectation is taken with respect to $\theta_{-j}\in\hat{\Theta}_{-j}$. Note that we are not worrying about the true distribution from bidder $j^{\prime}\neq j$ because when considering revenue from bidder $j$, we consider her expected transfers, and then integrate over $j$’s true expected distribution, $\theta_{j}\in\Theta^{*}_{j}$. Given that $v$ satisfies Assumption 5.1 and $\tilde{x}$ is orderly, we know that
$\mathds{E}$ $\left[\tilde{x}_{j}(\theta_{j},\theta_{-j})v(\theta_{j},\theta_{-j})\mid\theta_{j}=\theta_{j},\tilde{x}_{j}(\theta_{j},\theta_{-j})=1\right]$ (16)
$-C_{m}\cdot\left(\mathds{E}\left[v(\theta^{k_{m}},\theta_{-j})\mid\theta_{j}=\theta^{k_{m+1}},\tilde{x}_{j}(\theta^{k_{m}},\theta_{-j})=1\right]-\mathds{E}\left[v(\theta^{k_{m}},\theta_{-j})\mid\theta_{j}=\theta^{k_{m}},\tilde{x}_{j}(\theta^{k_{m}},\theta_{-j})=1\right]\right)$
also has the same properties (with respect to $\theta_{j}$) for any $m$ such that $\theta^{k_{m}}<\theta_{j}$ and constant $C_{m}$. Therefore, we can conclude that the necessary conditions for Lemma 5.6 are satisfied and apply it to conclude that
$\mathds{E}_{\theta\sim F^{\prime}}\left[T_{j}(\theta_{j};\hat{\Theta})-T_{j}(\theta_{j};\Theta^{*})\mid\theta\in\Theta^{*}\right]\geq\mathds{E}_{\theta\sim F}\left[T_{j}(\theta_{j};\hat{\Theta})-T_{j}(\theta_{j};\Theta^{*})\mid\theta\in\Theta^{*}\right].$
Thus, the auction is not shill-proof under $F^{\prime}$.
Common Values.
Towards contrapositive, suppose that the auction is not shill-proof (strong or weak, we will highlight where the proofs diverge) under $v$ and consider some $v^{\prime}\succeq_{\mathrm{Com}}v$. There exists some realization of $v$ under $v$ such that a shill bidder $i$ deviates at $\Theta$. Since $\tilde{x}$, $\mu$, and $\xi_{0}$ are all fixed, we know from Lemma 5.5 that under weak shill-proofness that the only way for expected revenue under weak shill-proofness to change is for the ex-interim expected value of some bidder $j\neq i$ to change. We can also apply Lemma B.3 to see that the same is true under strong shill-proofness. Thus, in order to prove that the auction is not shill-proof for either weak or strong shill-proofness under $F^{\prime}$, it is sufficient to show that for all $j\neq i$, if the expected transfers from $T_{j}$ from manipulation $\hat{\Theta}$ compared to truthful play $\Theta^{*}$ under $v$, then the same manipulation is profitable under $v^{\prime}$. This is immediate: since $v^{\prime}$ is point-wise higher than $v$ and more super-modular, then we can see that Equation 16 is larger under $v^{\prime}$ than $v$ no matter the constant and therefore the auction is not shill-proof.
Proof of Proposition 5.7
The proof that a strongly shill-proof auction must be pay-as-bid is the same as in Lemma 3.1 because that argument is ex-post and bidders’ values are weakly increasing in all types. To prove that the conjectured efficient allocation rule is efficient, observe that by the assumption that a bidder values her own signal more than that of others, the efficient ex-post allocation is such that $\tilde{x}_{i}=0$ for $i\notin\operatorname*{argmax}_{j}\left\{\theta_{j}\right\}$ and $\sum_{j}\tilde{x}_{j}(\theta)=1$ for all $\theta$. The assumption that the seller has 0 value for the good completes the proof that $\tilde{x}^{E}$ is the efficient allocation rule.
Recall that a shill-proof auction must be pay-as-bid, and observe that bids must be strictly increasing as a function of own signal in order to be incentive compatible. Next, because the bidder who has the highest ex-ante signal will have the highest ex-post valuation, under any allocation rule, the bidder with the highest signal who is allocated will have the highest bid. Therefore, to maximize transfers the optimal allocation rule is such that $\tilde{x}_{i}=0$ for $i\notin\operatorname*{argmax}_{j}\left\{\theta_{j}\right\}$. To conclude the proof, observe that by Lemma 5.5, the fact that changing the allocation rule conditional on only the maximum value will not change that bidder’s value, and by the standard arguments, the optimal allocation rule uses a reserve type.
Appendix E Supplemental Appendix